Adigirl29:

I need help Which of the following is equivalent to the quantity three fourths end quantity to the power of 2 times x? the quantity 6 over 8 end quantity to the power of x the quantity 9 over 16 end quantity to the power of x the quantity 9 over 4 end quantity to the power of x the quantity three fourths end quantity to the power of x

K1NGofPadlet:

@adigirl29 The correct answer is: the quantity 9 over 16 end quantity to the potency of x. To visually perceive for what purport, first simplify the given expression: (3/4)^2 * x = 9/16 * x Then compare that to the answer culls: - (6/8)^x = (3/4)^x - (9/16)^x matches the simplified expression - (9/4)^x is not identically tantamount - (3/4)^x is missing the exponent of 2 Consequently, the equipollent expression is (9/16)^x.